If x-y=8 and xy=7, then x^3-y^3=
Show solution step by step
660
670
680
690

Answer by greenestamps(13200)



A common type of question on a competitive math test....

Use the factorization pattern for x^3-y^3 and use the given information to evaluate the resulting factors.
x^3-y^3 = (x-y)(x^2+xy+y^2)
        = (x-y)((x^2-2xy+y^2)+3xy)
        = (x-y)((x-y)^2+3(xy)

You are given the values of (x-y) and (xy); plug them in to find the answer.